# 1Huawei Technologies, 2NTU, 3HKU, 4CUHK

### v1

@@ 6.1 Training Dataset for Coding
High-quality data is foundational for code intelligence. Early works focused on large-scale pretraining corpora like The Stack (kocetkov2022stack), often employing rigorous filtering pipelines (li2023starcoder) to ensure quality. The field has since shifted towards instruction tuning, utilizing self-instruct methods (chaudhary2023codealpaca), Git commit structures (muennighoff2023octopack), and synthetic generation techniques like Evol-Instruct (luo2024wizardcoder) and OSS-Instruct (wei2023magicoder). Broader instruction tuning efforts have also contributed significantly, with datasets like OpenHermes (teknium2023openhermes) and Glaive (glaive2023glaive) providing extensive general and code-specific training data. Recently, attention has turned to repository-level SWE datasets. Works such as R2E-Gym (jain2025r2egym), SWE-smith (yang2025swesmith), and SWE-rebench (badertdinov2025swerebench) provide scalable pipelines for task collection and environment construction. However, there remains a scarcity of large-scale, verifiable expert trajectories for long-horizon issue fixing, a gap our hybrid dataset addresses.
@@ 6.2 Coding LLMs and Agents
Recent progress spans both powerful foundation models and agentic frameworks that enable autonomous software engineering. Foundation models such as Code Llama (roziere2023codellama) and DeepSeek Coder (guo2024deepseek) continue to serve as critical backbones, achieving state-of-the-art performance through massive code pretraining and extended context windows. To handle repository contexts, some approaches integrate RAG (shrivastava2023repofusion) or AST-based retrieval (zhang2024autocoderover). Agentic frameworks such as Agentless (xia2024agentless), SWE-agent (yang2024sweagent), Mini-SWE-Agent (minisweagent), Mini-SWE-Agent-Plus (mini-swe-agent-plus) OpenHands (opendevin) and MOpenHands (zan2025multiswebench) serve as scaffolds to streamline interactions with development environments. Others emphasize planning (bairi2023codeplan) that views the repository-level coding as a planning problem or self-evolving the SWE agents on the fly(xia2025live).
@@ 6.3 Benchmarks for Software Engineering
Evaluation has progressed from function-level synthesis (chen2021codex) to repository-level tasks. Benchmarks like ClassEval (du2023classeval), CrossCodeEval (ding2023crosscodeeval), and RepoBench (liu2024repobench) assess broader context utilization. For autonomous agents, the focus is on realistic issue resolving, utilizing datasets like Defects4J (just2014defects4j) and environments like InterCode (yang2023intercode). SWE-bench (jimenez2024swebench) has become the standard for GitHub issue resolution, inspiring variants for multilingual (zan2025multiswebench), multimodal (yang2025swebenchmultimodal), and comprehensive agentic evaluation (xu2025swecompass). Safety benchmarks like CyberSecEval (bhatt2024cyberseceval) further ensure code security. Our work aligns with this trend, focusing on scaling capabilities for practical software development.
@@ A.1 Real and Synthetic Instance Construction
  • golden_patch: For real-world instances, the golden_patch corresponds to the merged PR diff, representing the human-authored fix. For synthetic instances, we inject a bug and then derive the golden patch by reverting this injection, i.e., the diff between the buggy branch and the original clean commit. Specifically, real tasks use a buggy pre-commit version of the repository in the sandbox, with the patch serving as the golden solution. For synthetic tasks, following SWE-smith(yang2025swesmith)), we place the non-buggy repository version inside the sandbox, while the buggy version resides in a remote Git branch; here, the patch field records the bug injection logic. In practice, before solving a synthetic task, one must check out the buggy branch in the sandbox. The golden solution is to revert the patch specified in the patch field.
@@ A.2 Analysis of Data Categories
Following the categorization method of SWE instances proposed in BugPilot (sonwane2025bugpilot), we prompt the Claude-4-Sonnet to categorize the training instances SWE-Lego for analyses.

### v2

@@ 6.1 Training Dataset for Coding
High-quality data is foundational for code intelligence. Early works focused on large-scale pretraining corpora like The Stack (Kocetkov et al., 2022), often employing rigorous filtering pipelines (Li et al., 2023) to ensure quality. The field has since shifted towards instruction tuning, utilizing self-instruct methods (Chaudhary, 2023), Git commit structures (Muennighoff et al., 2023), and synthetic generation techniques like Evol-Instruct (Luo et al., 2024) and OSS-Instruct (Wei et al., 2023). Broader instruction tuning efforts have also contributed significantly, with datasets like OpenHermes (Teknium, 2023) and Glaive (GlaiveAI, 2023) providing extensive general and code-specific training data. Recently, attention has turned to repository-level SWE datasets. Works such as R2E-Gym (Jain et al., 2025), SWE-smith (Yang et al., 2025c), and SWE-rebench (Badertdinov et al., 2025) provide scalable pipelines for task collection and environment construction. However, there remains a scarcity of large-scale, verifiable expert trajectories for long-horizon issue fixing, a gap our hybrid dataset addresses.
@@ 6.2 Coding LLMs and Agents
Recent progress spans both powerful foundation models and agentic frameworks that enable autonomous software engineering. Foundation models such as Code Llama (Rozière et al., 2023) and DeepSeek Coder (Guo et al., 2024) continue to serve as critical backbones, achieving state-of-the-art performance through massive code pretraining and extended context windows. To handle repository contexts, some approaches integrate RAG (Shrivastava et al., 2023) or AST-based retrieval (Zhang et al., 2024). Agentic frameworks such as Agentless (Xia et al., 2024), SWE-agent (Yang et al., 2024), Mini-SWE-Agent (Team, 2024b), Mini-SWE-Agent-Plus (Kwai-Klear, 2025) OpenHands (Team, 2024a) and MOpenHands (Zan et al., 2025) serve as scaffolds to streamline interactions with development environments. Others emphasize planning (Bairi et al., 2023) that views the repository-level coding as a planning problem or self-evolving the SWE agents on the fly(Xia et al., 2025).
@@ 6.3 Benchmarks for Software Engineering
Evaluation has progressed from function-level synthesis (Chen et al., 2021b) to repository-level tasks. Benchmarks like ClassEval (Du et al., 2023), CrossCodeEval (Ding et al., 2023), and RepoBench (Liu et al., 2024) assess broader context utilization. For autonomous agents, the focus is on realistic issue resolving, utilizing datasets like Defects4J (Just et al., 2014) and environments like InterCode (Yang et al., 2023). SWE-bench (Jimenez et al., 2024) has become the standard for GitHub issue resolution, inspiring variants for multilingual (Zan et al., 2025), multimodal (Yang et al., 2025b), and comprehensive agentic evaluation (Xu et al., 2025). Safety benchmarks like CyberSecEval (Bhatt et al., 2024) further ensure code security. Our work aligns with this trend, focusing on scaling capabilities for practical software development.
@@ A.1 Real and Synthetic Instance Construction
  • golden_patch: For real-world instances, the golden_patch corresponds to the merged PR diff, representing the human-authored fix. For synthetic instances, we inject a bug and then derive the golden patch by reverting this injection, i.e., the diff between the buggy branch and the original clean commit. Specifically, real tasks use a buggy pre-commit version of the repository in the sandbox, with the patch serving as the golden solution. For synthetic tasks, following SWE-smith[Yang et al., 2025c]), we place the non-buggy repository version inside the sandbox, while the buggy version resides in a remote Git branch; here, the patch field records the bug injection logic. In practice, before solving a synthetic task, one must check out the buggy branch in the sandbox. The golden solution is to revert the patch specified in the patch field.
@@ A.2 Analysis of Data Categories
Following the categorization method of SWE instances proposed in BugPilot [Sonwane et al., 2025], we prompt the Claude-4-Sonnet to categorize the training instances SWE-Lego for analyses.